The annual number of earthquakes registering at least 2.5 on the Richter Scale and having an epicenter within 40 miles of downtown Memphis follows a Poisson distribution with mean 6.5.

a. What is the probability that at least 9 such earthquakes will strike next year?

b. What is the probability that at least 25 such earthquakes will strike during the next 5 years?


1
Expert's answer
2021-02-24T06:50:38-0500

Solution:

a.

We need to find the probability of Poisson distribution wil "\lambda = 6.5"

Remembering Poisson formula and using calculator:


"P_n(\xi=m) = \frac{\lambda^m}{m!}e^{-\lambda}"


"P_n(\xi\ge9)=1- P_n(\xi=0)-P_n(\xi=1)-...-P_n(\xi=8)="

"=1 - 0.0015 - 0.00977-0.03176-0.06881-0.11182-0.14537-"

"-0.15748-0.14623-0.11882\approx0.20843"


b.

We can use the other Poisson distribution, for 5 years, where "\lambda=6.5*5=32.5"

Using inverted probability:


"P_n(\xi\ge25)= 1- P_n(\xi \le 24)= 1 - P_n(\xi=0) - P_n(\xi=1) - ...-P_n(\xi=24) ="

"= 1- 0.07536 =0.92464"


Answer:

a. "P_n(\xi\ge 9)=0.20843"

b."P_n(\xi\ge 25)=0.92464"